Max Powers, Vice President for National Lending, has instructed the computer programmer to use a 365-day year to calculate interest expense on the company's liabilities, which are loans they have taken out. He has also instructed the programmer to use a 360-day year to calculate interest revenue on assets, which are the loans they have given out to clients. 1. What are the financial ramifications of using a 365-day year opposed to a 360-day year on both the interest expense and interest revenue? 2. Which method should be used? 3. Can both methods be used? 4. Discuss whether Max is behaving in a professional manner. 5. Is there a method that banks normally use on loans?
Answer #1

We will discuss this problem with the help of an example :

Let says :

Loan Amount 500000
Interest rate 10% (Simply)
Annual Interest Will be 500000*10% = 50000pa

Concept of 360Days/Year Concept of 365Days/Year
Intt./Day : 50000/360 =138.88 Intt./Day : 50000/365 = 136.99
Annual Interest : 138.88*365 = 506910.00 (Round off) Annual Interest : 50001.00 (Round off)
Result : Better to Give Result : Better to Receive
Interest revenue on assets should be Recognized as per 360days/year concepts Interest expenses on Liability should be recorded as per 365days/year concept
Income should be booked Expenses should be recorded
Also Known as Bank Interest Method Called Stated Interest Method.

Answer : As you see its very clear with the example above, a rational way of policy for interest value. a professional will always go with the both the options.

1. Mr. Max Powers, VP for national lending instruct correctly and in his decision is rational. Financial ramification of using a 365-days year opposed to a 360-days year on both the interest expenses and interest revenue.
2. As National Lending is funding to clients it should follow 360days/year concept in case of interest income.

3. Yes, Both method can be used in the company however in case of OUTSIDE FUNDING always go with 360-days/year and INCOMING LOAN 365-Days/year should be follow.

4. with the example its clarify that Mr. Max is professional and his decision is rational.
5. Banks are always follows 360-days/year Method.